OpenStudy (anonymous):

A tank is full of water. Find the work W required to pump the water out of the spout. (Use 9.8 for g and 3.14 for π. If you enter your answer in scientific notation, round the decimal value to two decimal places. Use equivalent rounding if you do not enter your answer in scientific notation.) r=6.6, h=3.3

OpenStudy (anonymous):

it is spherical

OpenStudy (therealmeeeee):

W = gπr^2(integral {hdh } = gπr^2(h^2)/2 I'll leave it to you to plug in the proper values and round off the answer. In the future, it would be nice if you attach the proper unit to the various measurements. e.g. h=1.5 is it measured in inches, meters, yards ? I'll leave it to you to convert things to the proper units to make W = gπr^2(h^2)/2 correct.

OpenStudy (therealmeeeee):

hope this help click best response

OpenStudy (therealmeeeee):

do this help u @Jmnz

OpenStudy (anonymous):

thanks and i forgot to mention it is in meters

OpenStudy (therealmeeeee):

wattttttttttttttttt

OpenStudy (alekos):

Is this a cylindrical tank of water?

OpenStudy (anonymous):

no it is a spherical tank

OpenStudy (anonymous):

https://www.webassign.net/scalcet/6-4-022alt.gif

OpenStudy (therealmeeeee):

ohh

OpenStudy (alekos):

OK then the work required W is equal to the total mass of water multiplied by the total distance to pump that water from the bottom of the tank to the top of the spout (assuming there's a hose going down the spout to the bottom) So we have W = mgd where m= mass of water, g= 9.8 and d=r+h m = Volume x density of water m= (4/3)pi.r^3 x 1000 I'll let you do the rest

OpenStudy (anonymous):

ok thanks but i do not understand how to do the rest.

OpenStudy (alekos):

to get the work done W you just multiply mass x g x d the mass is (4/3) x pi x r^3 x 1000 g = 9.8 and d = r+h and you know r=3.3 and h=6.6 you follow now?

OpenStudy (anonymous):

yea i follow now... thank you

OpenStudy (alekos):

OK, good. the final answer will be in Joules